Bruno's Lunch Counter is expanding and expects operating cash flows of $28,500 a year for 6 years as a result. This expansion requires $95,800 in new fixed assets. These assets will be worthless at the end of the project. In addition, the project requires $7,000 of net working capital throughout the life of the project. What is the net present value of this expansion project at a required rate of return of 10 percent?
Answer #1

NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

C0 = 95800+7000 = 102800

CF1 to CF5 = 28500

CF6 = 28500+7000 (working capital recovery)

= 35500

NPV = -102800+ 28500/1.1^1 + 28500/1.1^2+ 28500/1.1^3+ 28500/1.1^4+ 28500/1.1^5+ 35500/1.1^6

=

25276.25

The NPV is $25276.25